# Neutrino flavor instabilities in a binary neutron star merger remnant: Roles of a long-lived hypermassive neutron star

### v1

@@ III.2.3 ELN crossings with a disk shadow
At the time snapshot of $t=0\,$s, ELN crossings appear in the equatorial region at a radius greater than $\sim 60\,$km (see the top left panel in Fig. 3). As shown below, the inner region of accretion disk with $\mu_{\nu}^{\rm eq}>0$ is responsible for them.
@@
One may wonder why ELN angular crossings are not observed in higher lattitudes (above the disk). This can be illustrated by Ray 4 and Ray 4-f in Fig. 9. In the vicinity of energy sphere of $\nu_{e}$, ELN is certainly positive which can generate ELN angular crossings. In fact, ELN angular crossings appear at the edge of optically thick region with $\mu_{\nu}^{\rm eq}>0$ (see top left panel in Fig. 3), whose mechanism is essentially the same as the one with $\nu_{e}$ diffusion (see Sec. III.2.2). As already mentioned above, however, neutron rich environments above the disk attenuate $\nu_{e}$ due to charged-current reactions ($\nu_{e}$ absorption by neutrons), and $\bar{\nu}_{e}$ emitted from HMNS is more likely to scatter off nucleons. Consequently, ELN eventually becomes negative as neutrinos propagate towards higher latitudes (see the color transition from Ray 4 to Ray 4-f in Fig. 9), causing disappearance of ELN angular crossings.
@@
It should be emphasized that HMNS also plays a crucial role on the mechanism. In fact, if HMNS has already collapsed into a BH, there would be no angular regions with negative ELN at the corresponding region. Let us also point out that the emergence of $\mu_{\nu}^{\rm eq}>0$ region plays an important role similar to other mechanisms. However, this is merely a necessary condition for appearance of this type of ELN angular crossings. If the positive $\mu_{\nu}^{\rm eq}$ region is surrounded by optically thick medium with negative $\mu_{\nu}^{\rm eq}$, neutrinos emitted from the disk would be dominated by $\bar{\nu}_{e}$, erasing ELN angular crossings. In addition to this, the disk needs to be optically thick to block neutrinos from HMNS, indicating that the ELN crossing does not appear in the late phase after the disk becomes transparent. In fact, this type of crossings is no longer observed after 0.20.20.2\,0.2s in our model.
@@ III.2.4 ELN crossings by $\nu_{e}$ contamination
After $\sim 0.2\,$s, an optically thick accretion disk with $\mu_{\nu}^{\rm eq}>0$ disappears, while regions with $\mu_{\nu}^{\rm eq}\sim 0$ emerge above the disk. Before delving into ELN angular crossings, we briefly discuss them, since they play an important role in the generation of such crossings.
@@
Let us consider ELN angular crossings associated with the band like structure of $\mu_{\nu}^{\rm eq}\gtrsim 0$. Fig. 12 presents a schematic illustration of the underlying mechanism. This figure is sketched to represent the case with $t=0.6\,$s (see the bottom left panels in Figs. 1 and 3) but the mechanism is essentially the same for other time snapshots ($t=0.8,$ and 1.31.31.3\,1.3s). As already pointed out in previous subsections, $\bar{\nu}_{e}$ emission from the surface of HMNS dominates over $\nu_{e}$, and the former undergoes more frequent scatterings with nucleons than the latter. This makes the ELN angular distributions be negative in the full angles. On the other hand, $\nu_{e}$ emission dominates over $\bar{\nu}_{e}$ in the band-like flow with $\mu_{\nu}^{\rm eq}\gtrsim 0$, which potentially changes the sign of ELN. In the radially forward directions, the intensity of $\bar{\nu}_{e}$ is strong (since they are emitted from the vicinity of HMNS; see Ray 2 in Fig. 12), implying that ELN becomes negative. In the backward directions, $\bar{\nu}_{e}$ is still more abundant than $\nu_{e}$ (see Ray 2b). This is attributed to the angular dependence of nucleon scatterings, which favors backward scatterings. Although $\bar{\nu}_{e}$ intensity is much lower than the forward direction, it still leads to negative ELN. Let us, then, consider the neutrinos coming from the disk region. In the equatorial region, $\bar{\nu}_{e}$ also dominates. This is due to stronger charged current reaction for $\bar{\nu}_{e}$ and more frequent $\bar{\nu}_{e}$ scatterings than those for $\nu_{e}$. As a result, more $\bar{\nu}_{e}$ can escape vertically from the equatorial region, that also makes ELN negative in this direction (see Ray 1s). However, neutrinos which propagate from higher latitude to the equatorial plane have a distinct behavior. In polar regions where matter density is very low, neutrino scatterings are less active (see Ray 3 and 3s). As a result, $\nu_{e}$ emission in the band-like flow with $\mu_{\nu}^{\rm eq}\gtrsim 0$ can exceed the scattered $\bar{\nu}_{e}$, resulting in a positive ELN. In Fig. 13, we show that the ELN becomes positive in the angular region of $\phi_{\nu}\sim 0(2\pi)$ and $\cos{\theta_{\nu}}\lesssim 0$, which supports our interpretation.
@@
It is worth mentioning that this type of ELN crossings tends to vanish in the equatorial region, which is illustrated as Ray 3s + Ray 1s in Fig. 12. In this region, the neutron abundance is higher than in higher latitudes, leading to many scattered $\bar{\nu}_{e}$ and local $\bar{\nu}_{e}$ emission through a charged current process. As a result, ELN becomes negative for Ray 3s + Ray 1s. This accounts for no ELN angular crossings in the equatorial region at $t=0.6\,$s time snapshot. It should be mentioned that the competition between $\bar{\nu}_{e}$ and $\nu_{e}$ hinges on the matter profile. The scattered neutrinos decrease with time due to lower neutrino emission from HMNS and low matter density in the disk (which also decrease with time). Consequently, $\bar{\nu}_{e}$ cannot overcome $\nu_{e}$ for Ray 3s + Ray 1s in Fig. 12, suggesting that ELN crossings can be observed in the equatorial region at later phases. This accounts for the spatial distributions for ELN angular crossings for later time snapshots of $t=0.8\,$s and 1.31.31.3\,1.3s (see Fig. 3).
@@
Let us make another argument to explain no ELN angular crossings at $\gtrsim 20\,$km for the time snapshots of $t=0.2\,$s and 0.40.40.4\,0.4s, despite the fact that there is a band-like structure of $\mu_{\nu}^{\rm eq}\gtrsim 0$. During these relatively early post-merger phases, $\bar{\nu}_{e}$ emission from HMNS is still strong, and the matter density in the polar region is also higher than in the late phase. As a result, the scattered neutrinos from the polar region (Ray 3s in Fig. 12) is abundant, which increase $\bar{\nu}_{e}$ intensity. As a result, $\nu_{e}$ emission in the region of $\mu_{\nu}^{\rm eq}\gtrsim 0$ cannot dominate over $\bar{\nu}_{e}$, suppressing ELN angular crossings.
Here is a short summary of this subsection regarding detailed analyses of FFI. The appearance of $\mu_{\nu}^{\rm eq}\gtrsim 0$ underlies all mechanism to generate ELN angular crossings. This suggests that, although the robust assessment of ELN angular crossings requires multi-angle neutrino transport simulations, $\mu_{\nu}^{\rm eq}$ distributions, which can be estimated from only thermodynamical quantities, can serve as a useful proxy for identifying occurrences of FFI in BNSM simulations. It should also be emphasized that intense $\bar{\nu}_{e}$ emission from HMNS leads to distinct crossing generation mechanisms from those in BH accretion disk systems.
@@ III.3.2 Stable configuration of CFI
Although the growth rate of CFI is much lower than FFI, the present study shows that CFI plays a complementary role on flavor conversions. For instances, the CFI structure is more persistent than FFI for $t\lesssim 1\,$s. In fact, the spatial region of CFI tends to be wider than FFI, and CFI can occur in time snapshots of $t=0.2\,$s and 0.40.40.4\,0.4s but FFI does not occur at outer radii ($\gtrsim 20\,$km) at these time snapshots. This result suggests that there are some phases which CFI must dominate over FFI.
@@ III.3.3 Disappearance of CFI in the late phase
In Fig. 16, we show the species-dependent neutrino number luminosity measured at the outer boundary as a function of a polar angle. For comparison, we display them for two different time snapshots: $t=0.2\,$s (left) and 1.31.31.3\,1.3s (right). As can be seen in this figure, the number luminosity of $\nu_{x}$ exceeds $\nu_{e}$ for $t=1.3\,$s, while $\bar{\nu}_{e}$ is the highest number flux. This suggests that the sign of $g$ and $\bar{g}$ (see Eq. (7)) is negative and positive, respectively. This leads to $|G|<|A|$ (see Eq. (6)), which makes the system stable with respect to CFI (see Eqs. (4) and (5)). Our finding suggests that the relative number densities of different neutrino species is a good diagnostic to assess occurrences of CFI.
@@ IV Summary and conclusion
    In our models, four different mechanisms operate in generating ELN angular crossings (see Sec. III.2). In each case, occurrences of ELN angular crossings are associated with regions with $\mu_{\nu}^{\rm eq}>0$. This is attributed to the fact that the overall radiation field is dominated by $\bar{\nu}_{e}$ due to the presence of HMNS. In such environments, $\nu_{e}$ supply by charged-current reactions in regions with $\mu_{\nu}^{\rm eq}>0$ can make ELN positive in along certain neutrino trajectories, which is indicative of appearance of ELN angular crossings. We refer readers to Sec. III.2 for more details. We note that three schematic pictures (Figs. 79, and 12) help readers to understand complex generation mechanisms graphically.
@@ Appendix A Fluid profiles
We provide matter profiles used in this study, which are obtained by numerical relativity simulations in [60] (DD2-135M model). We note that our previous work in [74] focused on the phase with $0<t<135\,$ms, while the present study extends over a longer timescale ($>1\,$s). Figure 17 portrays spatial distributions of baryon mass density at different time snapshots. We also provide distributions of matter temperature in Fig. 18. In analysis of flavor instabilities, these figures can be used as a reference displaying the overall trend of time evolution of the accretion disk. See also Figs. 1 and 2 for chemical potential of $\nu_{e}$ and $Y_{e}$, respectively.
@@ Appendix B Grid structure and resolution study in Boltzmann neutrino transport simulations
As mentioned in Sec. II.2, high spatial resolutions are necessary in modeling of neutrino transport in the vicinity of HMNS, where matter profiles are substantially changed over the short distance. We note that HMNS gradually shrinks with time due to neutrino cooling, and consequently the density gradient around the envelope becomes sharper. This indicates that the required spatial resolution needs to be higher in the later phase of BNSM remnant. According to our resolution study, insufficient spatial resolutions results in the overestimation of neutrino flux, and the least required resolution around the envelop of HMNS at $T=1.3\,$s (the final time snapshot in our simulation) is $\sim 20\,$m to accurately model the neutrino transport.
On the other hand, such a fine resolution is not required inside HMNS and in the outer envelope, where matter gradient is more moderate. We, hence, adopt a grid structure tailored to the problem. Figure 19 displays radial grid width as a function of radius for three different cases. The black dotted curve represents the case used in our previous study [74], while the red solid curve corresponds to the one adopted in the present study. As depicted, the new grid structure has a $\sim 20\,$m resolution at $r\sim 10\,$km where the density gradient is the steepest. On the other hand, the radial grid width at the center is roughly ten times larger than the finest resolution, and it increases logarithmically with radius at $r\gtrsim 15\,$km. This gridding strategy efficiently reduces the number of grid points while maintaining high spatial resolutions in the outer region of the HMNS.

### v2

@@ III.2.3 ELN crossings with a disk shadow
At the time snapshot of $t=0\,$s, ELN crossings appear in the equatorial region at a radius greater than 60$\sim$ 60\,∼ 60km (see the top left panel in Fig. 3). As shown below, the inner region of accretion disk with $\mu_{\nu}^{\rm eq}>0$ is responsible for them.
@@
One may wonder why ELN angular crossings are not observed in higher latitudes (above the disk). This can be illustrated by Ray 4 and Ray 4-f in Fig. 9. In the vicinity of energy sphere of $\nu_{e}$, ELN is certainly positive which can generate ELN angular crossings. In fact, ELN angular crossings appear at the edge of optically thick region with $\mu_{\nu}^{\rm eq}>0$ (see top left panel in Fig. 3), whose mechanism is essentially the same as the one with $\nu_{e}$ diffusion (see Sec. III.2.2). As already mentioned above, however, neutron rich environments above the disk attenuate $\nu_{e}$ due to charged-current reactions ($\nu_{e}$ absorption by neutrons), and $\bar{\nu}_{e}$ emitted from HMNS is more likely to scatter off nucleons. Consequently, ELN eventually becomes negative as neutrinos propagate towards higher latitudes (see the color transition from Ray 4 to Ray 4-f in Fig. 9), causing disappearance of ELN angular crossings.
@@
It should be emphasized that HMNS also plays a crucial role on the mechanism. In fact, if HMNS has already collapsed into a BH, there would be no angular regions with negative ELN at the corresponding region. Let us also point out that the emergence of $\mu_{\nu}^{\rm eq}>0$ region plays an important role similar to other mechanisms. However, this is merely a necessary condition for appearance of this type of ELN angular crossings. Similar to the argument in Sec. III.2.2, the excess of $\nu_{e}$ compared to $\bar{\nu}_{e}$ in low energy (10$\lesssim$ 10\,≲ 10MeV) regions needs to be sufficiently high to counteract the negative ELN in high energy regions. This condition is not trivial and hinges on the matter distribution around the disk. We also note that the disk needs to be optically thick to block neutrinos from HMNS for the shadow mechanism, indicating that the ELN crossing does not appear in the late phase after the disk becomes transparent. In fact, this type of crossings is no longer observed after 0.20.2\,0.2s in our model.
@@ III.2.4 ELN crossings by $\nu_{e}$ contamination
After 0.2$\sim$ 0.2\,∼ 0.2s, an optically thick accretion disk with $\mu_{\nu}^{\rm eq}>0$ disappears, while regions with $\mu_{\nu}^{\rm eq}\sim 0$ emerge above the disk. Before delving into ELN angular crossings, we briefly discuss them, since they play an important role in the generation of such crossings.
@@
Let us consider ELN angular crossings associated with the band like structure of $\mu_{\nu}^{\rm eq}\gtrsim 0$. Fig. 12 presents a schematic illustration of the underlying mechanism. This figure is sketched to represent the case with $t=0.6\,$s (see the bottom left panels in Figs. 1 and 3) but the mechanism is essentially the same for other time snapshots ($t=0.8,$ and 1.31.3\,1.3s). As already pointed out in previous subsections, $\bar{\nu}_{e}$ emission from the surface of HMNS dominates over $\nu_{e}$, and the former undergoes more frequent scatterings with nucleons than the latter. This makes the ELN angular distributions be negative in the full angles. On the other hand, $\nu_{e}$ emission dominates over $\bar{\nu}_{e}$ in the band-like flow with $\mu_{\nu}^{\rm eq}\gtrsim 0$, which potentially changes the sign of ELN. In the radially forward directions, the intensity of $\bar{\nu}_{e}$ is strong (since they are emitted from the vicinity of HMNS; see Ray 2 in Fig. 12), implying that ELN becomes negative. In the backward directions, $\bar{\nu}_{e}$ is still more abundant than $\nu_{e}$ (see Ray 2b). This is attributed to the angular dependence of nucleon scatterings, which favors backward scatterings. Although $\bar{\nu}_{e}$ intensity is much lower than the forward direction, it still leads to negative ELN. Let us, then, consider the neutrinos coming from the disk region. In the equatorial region, $\bar{\nu}_{e}$ also dominates. This is due to stronger charged current reaction for $\bar{\nu}_{e}$ and more frequent $\bar{\nu}_{e}$ scatterings than those for $\nu_{e}$. As a result, more $\bar{\nu}_{e}$ can escape vertically from the equatorial region, that also makes ELN negative in this direction (see Ray 1s). However, neutrinos which propagate from higher latitude to the equatorial plane have a distinct behavior. In polar regions where matter density is very low, neutrino scatterings are less active (see Ray 3 and 3s). As a result, $\nu_{e}$ emission in the band-like flow with $\mu_{\nu}^{\rm eq}\gtrsim 0$ can exceed the scattered $\bar{\nu}_{e}$, resulting in a positive ELN. In Fig. 13, we show that the ELN becomes positive in the angular region of $\phi_{\nu}\sim 0(2\pi)$ and $\cos{\theta_{\nu}}\lesssim 0$, which supports our interpretation.
@@
It is worth mentioning that this type of ELN crossings tends to vanish in the equatorial region, which is illustrated as Ray 3s + Ray 1s in Fig. 12. In this region, the neutron abundance is higher than in higher latitudes, leading to many scattered $\bar{\nu}_{e}$ and local $\bar{\nu}_{e}$ emission through a charged current process. As a result, ELN becomes negative for Ray 3s + Ray 1s. This accounts for no ELN angular crossings in the equatorial region at $t=0.6\,$s time snapshot. It should be mentioned that the competition between $\bar{\nu}_{e}$ and $\nu_{e}$ hinges on the matter profile. The scattered neutrinos decrease with time due to lower neutrino emission from HMNS and low matter density in the disk (which also decrease with time). Consequently, $\bar{\nu}_{e}$ cannot overcome $\nu_{e}$ for Ray 3s + Ray 1s in Fig. 12, suggesting that ELN crossings can be observed in the equatorial region at later phases. This accounts for the spatial distributions for ELN angular crossings for later time snapshots of $t=0.8\,$s and 1.31.3\,1.3s (see Fig. 3).
@@
Let us make another argument to explain no ELN angular crossings at 20$\gtrsim$ 20\,≳ 20km for the time snapshots of $t=0.2\,$s and 0.40.4\,0.4s, despite the fact that there is a band-like structure of $\mu_{\nu}^{\rm eq}\gtrsim 0$. During these relatively early post-merger phases, $\bar{\nu}_{e}$ emission from HMNS is still strong, and the matter density in the polar region is also higher than in the late phase. As a result, the scattered neutrinos from the polar region (Ray 3s in Fig. 12) is abundant, which increase $\bar{\nu}_{e}$ intensity. As a result, $\nu_{e}$ emission in the region of $\mu_{\nu}^{\rm eq}\gtrsim 0$ cannot dominate over $\bar{\nu}_{e}$, suppressing ELN angular crossings.
Here is a short summary of this subsection regarding detailed analyses of FFI. The appearance of $\mu_{\nu}^{\rm eq}\gtrsim 0$ underlies all mechanism to generate ELN angular crossings. However, it is important to emphasize that this condition is not sufficient on its own to guarantee occurrences of FFI as discussed above, and multi-angle neutrino transport simulations are necessary for the reliable assessment. We also illustrated in this section that intense $\bar{\nu}_{e}$ emission from HMNS leads to distinct crossing generation mechanisms from those in BH accretion disk systems.
@@ III.3.2 Stable configuration of CFI
Although the growth rate of CFI is much lower than FFI, the present study shows that CFI plays a complementary role on flavor conversions. For instances, the CFI structure is more persistent than FFI for $t\lesssim 1\,$s. In fact, the spatial region of CFI tends to be wider than FFI, and CFI can occur in time snapshots of $t=0.2\,$s and 0.40.4\,0.4s but FFI does not occur at outer radii (20$\gtrsim$ 20\,≳ 20km) at these time snapshots. This result suggests that there are some phases which CFI must dominate over FFI.
@@ III.3.3 Disappearance of CFI in the late phase
[…]
@@ IV Summary and conclusion
    In our models, four different mechanisms operate in generating ELN angular crossings (see Sec. III.2). In each case, occurrences of ELN angular crossings are associated with regions with $\mu_{\nu}^{\rm eq}>0$. This is attributed to the fact that the overall radiation field is dominated by $\bar{\nu}_{e}$ due to the presence of HMNS. In such environments, $\nu_{e}$ supply by charged-current reactions in regions with $\mu_{\nu}^{\rm eq}>0$ can make ELN positive in along certain neutrino trajectories, which is indicative of appearance of ELN angular crossings. However, it is important to emphasize that ELN crossings do not occur unless the emission of $\nu_{e}$ is strong enough to exceed $\bar{\nu}_{e}$, indicating that the appearance of $\mu_{\nu}^{\rm eq}>0$ is merely a necessary condition for FFI (see Sec. III.2 for more details). We also note that three schematic pictures (Figs. 79, and 12) help readers to understand complex generation mechanisms graphically.
@@ Appendix A Fluid profiles
We provide matter profiles used in this study, which are obtained by numerical relativity simulations in [60] (DD2-135M model). We note that our previous work in [74] focused on the phase with $0<t<135\,$ms, while the present study extends over a longer timescale (>1>1\,> 1s). Figure 17 portrays spatial distributions of baryon mass density at different time snapshots. We also provide distributions of matter temperature in Fig. 18. In analysis of flavor instabilities, these figures can be used as a reference displaying the overall trend of time evolution of the accretion disk. See also Figs. 1 and 2 for chemical potential of $\nu_{e}$ and $Y_{e}$, respectively.
@@ Appendix B Grid structure and resolution study in Boltzmann neutrino transport simulations
As mentioned in Sec. II.2, high spatial resolutions are necessary in modeling of neutrino transport in the vicinity of HMNS, where matter profiles are substantially changed over the short distance. We note that HMNS gradually shrinks with time due to neutrino cooling, and consequently the density gradient around the envelope becomes sharper. This indicates that the required spatial resolution needs to be higher in the later phase of BNSM remnant. According to our resolution study, insufficient spatial resolutions results in the overestimation of neutrino flux, and the least required resolution around the envelop of HMNS at $T=1.3\,$s (the final time snapshot in our simulation) is 20$\sim$ 20\,∼ 20m to accurately model the neutrino transport.
On the other hand, such a fine resolution is not required inside HMNS and in the outer envelope, where matter gradient is more moderate. We, hence, adopt a grid structure tailored to the problem. Figure 19 displays radial grid width as a function of radius for three different cases. The black dotted curve represents the case used in our previous study [74], while the red solid curve corresponds to the one adopted in the present study. As depicted, the new grid structure has a 20$\sim$ 20\,∼ 20m resolution at $r\sim 10\,$km where the density gradient is the steepest. On the other hand, the radial grid width at the center is roughly ten times larger than the finest resolution, and it increases logarithmically with radius at $r\gtrsim 15\,$km. This gridding strategy efficiently reduces the number of grid points while maintaining high spatial resolutions in the outer region of the HMNS.